OpenStudy (help_people):

Hi this question I am not sure about, I am debating so may you please help me confirm the correct answer (Algebra)

OpenStudy (help_people):

The table below shows four systems of equations: System 1 System 2 System 3 System 4 4x − 5y = 2 3x − y = 8 4x − 5y = 2 10x − 7y = 18 4x − 5y = 2 3x − 8y = 4 4x − 5y = 2 10x + 3y = 15 Which pair of systems will have the same solution? System 1 and system 2, because the second equation in system 2 is obtained by adding the first equation in system 1 to two times the second equation in system 1 System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to two times the second equation in system 2 System 1 and system 2, because the second equation in system 2 is obtained by adding the first equation in system 1 to three times the second equation in system 1 System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to three times the second equation in system 2

OpenStudy (help_people):



OpenStudy (help_people):



OpenStudy (help_people):

@Nnesha

OpenStudy (anonymous):

15

OpenStudy (help_people):

? @0404diego

OpenStudy (anonymous):

hi

OpenStudy (help_people):

@mathstudent55

OpenStudy (help_people):

hi 04